OpenStudy (anonymous):

Express 2,558 in scientific notation. A 2.558 x 10^4 B 2.558 x 10^-4 C 2.558 x 10^3 D 2.558 x 10^-3

OpenStudy (anonymous):

I forgot how to do scientific notation

OpenStudy (anonymous):

2.558*10^4=25580 2.558*10^(-4)=0.0002558 2.558*10^3=2558 2.558*10^(-3)=0.002558